### v3

# GTPO: Trajectory-Based Policy Optimization in Large Language Models

Abstract
Policy-based optimizations are widely adopted today for the training and alignment of language models, where one of the most recent and effective approaches is Group-relative Policy Optimization (GRPO). In this paper, we reveal and analyze two major limitations of GRPO: (i) tokens frequently appear in completions with both positive and negative rewards, leading to conflicting gradient updates that can reduce their output probability, even though can be essential for maintaining proper structure; (ii) negatively rewarded completions may penalize confident responses and shift model decisions toward unlikely tokens, progressively flattening the output distribution and degrading learning. To address these issues and provide a more stable and effective policy optimization strategy, we introduce GTPO (Group-relative Trajectory-based Policy Optimization), which identifies conflict tokens, tokens appearing in the same position across completions with opposite rewards and protects them by skipping negative updates, while amplifying positive ones. To further prevent policy collapse, GTPO filters out completions whose entropy exceeds a provable threshold. Unlike GRPO, GTPO does not rely on KL-divergence regularization, eliminating the need for a reference model during training, while still ensuring greater training stability and improved performance, validated through multiple experiments on GSM8K, MATH and AIME 2024 benchmarks. The code is available on Github at this link [^1] http://bit.ly/4oV8c49.
1 Introduction
Recent advancements in the training and alignment of large language model (LLM) have led to the adoption of policy-based optimization techniques to encourage LLMs on matching desired human preferences. Prominent approaches such as DPO (Rafailov et al. 2023), RLHF (Ouyang et al. 2022), and RFT (Yuan et al. 2023) were among the first to incorporate human feedback to guide model responses during training, acting as Reinforcement Learning (RL) techniques (Shao et al. 2024). More recently, a significant advancement in RL was introduced with Group-Relative Policy Optimization (GRPO) (Shao et al. 2024), which represents a variant of PPO (Schulman et al. 2017) that eliminates the need for a specific critic model by estimating the baseline directly from grouped reward values. In GRPO, the LLM acts as a policy that generates step-by-step reasoning and receives deterministic rewards based on the aggregated correctness and formatting of multiple tentative answers (referred to as completions) inferred for each question.
Despite the improvements introduced by GRPO, there are two key issues identified and analyzed in this work: (i) undesired gradient conflicts affect potentially corrects tokens shared across multiple completions of the same group that receive both positive and negative advantages, and (ii) a policy collapse phenomenon, defined as a degradation in LLM performance after a certain number of training steps, in which negatively rewarded completions destabilize the training process. More specifically, we observe that the first issue primarily impacts formatting tokens, which are essential to ensure proper answer structure and style.
To address these issues, this paper proposes Group-relative Trajectory-based Policy Optimization (GTPO), which treats the sequence of generated tokens (i.e., the completion) as a trajectory of decisions taken by the LLM policy. The core idea behind GTPO is to prevent undesired divergence among trajectories within the same group, thus keeping stability, while enhancing the rewards. To do so, GTPO applies a conflict-aware gradient correction mechanism that mitigates gradient conflicts on shared tokens, particularly in the initial and final parts of completions, thus preserving formatting consistency across trajectories. Furthermore, we demonstrate that while KL divergence of GRPO often reacts too slowly in preventing policy collapse, monitoring the entropy in LLM outputs offers a clearer signal of policy instability. Building on this, entropy-based regularization terms are applied in GTPO to control the exploration of trajectories in the same group.
These two components introduced by GTPO prevents the penalization of important formatting tokens and mitigates policy collapse, improving both the structure and accuracy of generated completions. Notably, since GTPO no longer relies on KL divergence, it does not require a reference model during training, making the process more lightweight and faster. In summary, the contributions of the work are:
  • We identify and analyze two critical issues in GRPO, such as gradient conflicts on tokens shared between completions of the same groups; and a policy collapse phenomenon, highlighting the limitations of the KL term;
  • We introduce GTPO, a new policy-based optimization that addresses the previous issues via conflict-aware gradient corrections and entropy-based regularizations;
  • We conduct extensive experiments and ablations studies to validate the effectiveness of GTPO, demonstrating more stable training and improved performance on both in-distribution and out-of-distribution benchmarks, including GSM8K, MATH, and AIME2024, using LLaMA-8B and Qwen 2.5-3B.
The remainder of the paper is organized as follows: we first discuss related work and introduce important preliminaries. We then analyze and discuss issues identified in GRPO, followed by the description of GTPO. Finally, the paper presents the experimental results and states conclusions.
2 Related Work
Reinforcement Learning in LLMs. RL has been widely adopted in decision-making tasks (Mnih et al. 2016, 2015; Berner et al. 2019), and nowadays is increasingly applied to the alignment and fine-tuning of LLMs. A prominent approach is RL from Human Feedback (RLHF), introduced in InstructGPT (Ouyang et al. 2022), and further developed by Anthropic (Bai et al. 2022). RLHF has become central to the training pipelines of state-of-the-art LLMs such as Claude 3 (Anthropic 2024), Gemini (Anil et al. 2023), and GPT-4 (OpenAI 2023). It typically includes supervised fine-tuning, a reward model, and the adoption of Proximal Policy Optimization (PPO) (Schulman et al. 2017). In this settings, PPO improves training stability by constraining updates through a clipped surrogate objective, offering a more practical alternative to Trust Region Policy Optimization (TRPO) (Schulman et al. 2015). However, it remains sensitive to reward scaling and can suffer from training instability (Wang et al. 2019; Garg et al. 2021; Moalla et al. 2024), requiring multiple refinements (Huang et al. 2022). Therefore, multiple variations have been proposed over the years, such as TRGPPO (Wang et al. 2019), alphaPPO (Xu et al. 2023), and PPO-ALR (Jia et al. 2024).
Advancements and Limitations in GRPO. To overcome the need for a critic model, Deepseek introduced GRPO (Shao et al. 2024; Guo et al. 2025), an approach that, given a question, compares multiple responses (completions) to derive relative rewards. GRPO demonstrates state-of-the-art performance on math benchmarks and achieves human-like alignment without relying on explicit manual feedback or critic networks (Li et al. 2025). Despite its promise, potential limitations of GRPO have been recently emerged, as bias effects (He, Fried, and Welleck 2025), gradient imbalance (Yang et al. 2025b), which lead to undertraining of rare yet informative tokens (Liu et al. 2025), and degradations (or even collapse) of model performance like in PPO (Dohare, Lan, and Mahmood 2023).
Building upon previous analyses of GRPO training behaviors, we deepen the study of token-level updates across completions of the same group, revealing conflicts in shared tokens. Furthermore, we extend the understanding of policy collapse, showing that KL divergence is limited in addressing this issue, whereas entropy-based analysis provides clearer signals (Cui et al. 2025). These insights motivate the design of GTPO, which effectively improves stability and performance during both training and evaluation.
3 Preliminaries
In GRPO, the LLM, acting as a policy, generates during training multiple completions (responses) per prompt $q$, denoted as $\{o_{1},o_{2},\ldots,o_{G}\}$, where $G$ is the number of completions generated, and computing rewards relatively to each rather than absolutely. Each output is structured with special tags for reasoning and answer segments and allows the reward to be calculated as an aggregate of formatting and factual correctness. The goal is to maximize the objective:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q,\{{o}_{i}\}}\left[\frac{1}{G}\sum_{i=1}^{G}\bar{\mathcal{C}}_{i}-\beta\cdot D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}})\right]$ (1)
Here, the expectation is taken over prompts $q\sim\mathbb{P}(Q)$ and completions $\{o_{i}\}\sim\pi_{\theta_{\text{old}}}$, where $\pi_{\theta_{\text{old}}}$ is the behavior policy. Each completion $o_{i}$ has length $|o_{i}|$. The term $\bar{\mathcal{C}}_{i}$ denotes the average clipped advantage over its tokens: $\bar{\mathcal{C}}_{i}=\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\mathcal{C}_{i,t}.$ The clipped advantage for the token at position $t$ in the completion $o_{i}$, is defined as: $\mathcal{C}_{i,t}=\min\left(\frac{\pi_{\theta}(o_{i,t}\mid s_{i,t})}{\pi_{\theta_{\text{old}}}({o}_{i,t}\mid s_{i,t})}\cdot\hat{A}_{i},\ \text{clip}\right),$ where $\hat{A}_{i}=\frac{R_{i}-\bar{R}}{\mathrm{std}(R)}$ is the normalized scalar advantage assigned to the $i$-th completion, and $s_{i,t}=(q,o_{i,<t})$ denotes the sequence of previously generated tokens up to position $t$. Specifically $R_{i}$ represents the reward for completion $i$, which is a sum of sub-rewards based on the correctness of the answer and its formatting style (Formatting Reward). The term $\pi_{\theta}(o_{i,t}|s_{i,t})=\mathrm{softmax}(f_{\theta})\in\mathbb{R}^{V},$ represents the probability distribution over the output logits $f_{\theta}$, where $f^{j}_{\theta}$ is the logit at position $j$, given the context $s_{i,t}$, and $V$ is the vocabulary size. Finally, the KL divergence term in Eq.(1) penalizes deviation from a reference policy $\pi_{\mathrm{ref}}$, scaled by a factor $\beta$. Due to computational cost and training time, one iteration is adopted in Eq. 1 (Simoni et al. 2025), where $\pi_{\theta}=\pi_{\theta_{\text{old}}}$, making the likelihood ratio 1 and the clipping unnecessary, i.e., $\mathcal{C}_{i,t}=\hat{A}_{i}$. The gradient of this simplified objective is:
$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{A}_{i}}{|{o}_{i}|}\sum_{t=1}^{|{o}_{i}|}g_{i,t}-\beta\cdot\nabla_{\theta}\mathrm{D}_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}}),$ (2)
where $g_{i,t}:=\nabla_{\theta}\log\pi_{\theta}({o}_{i,t}|s_{i,t})$ denotes the token-level policy gradient, which can be also expressed as:
$g_{i,t}=\Big{(}\nabla_{\theta}f_{\theta}^{j}(1-\pi_{\theta}^{j})-\sum_{k\neq j}\pi_{\theta}^{k}\nabla_{\theta}f_{\theta}^{k}\Big{)},$
with $j$ is the index of the logit with the highest probability $\pi_{\theta}$. Intuitively from Eq.(2), a positive advantage ($\hat{A}_{i}>0$) boosts the logit of the selected token proportionally to $1-\pi^{j}$, while reducing the logits of all others. Converselly, a negative advantage inverts this behavior, penalizing the token and increasing alternatives. Full derivations are in Append.
4 GRPO Issues
This section highlights two major limitations of GRPO: Token-level penalization and policy collapse.
4.1 Token-level Penalization
We show that GRPO tends to have negative effects on the updates of certain tokens that are essential for maintaining the structure and interpretability of completions.
Prefix Tokens Penalization. Given a prompt $q$, the model generates a set of completions $\{o_{1},o_{2},\ldots,o_{G}\}$, which may begin with the same sequence of tokens and then diverge at a specific point, a token that acts as a crossroads. This behavior is typical in instruction-tuned models, where early tokens tend to be similar across completions (Wang et al. 2024). However, not all completions in the group $G$ necessarily share the same prefix. In practice, we often observe multiple distinct prefixes, each shared by a subset of the completions. To model this, we partition the $G$ completions into $K$ disjoint groups $\mathcal{G}_{1},\ldots,\mathcal{G}_{K}$ so that all completions in a group $\mathcal{G}_{k}$ have a common prefix $S_{\text{pfx}}^{(k)}$. If all $G$ completions share the same prefix, then $K=1$. Tokens in $S_{\text{pfx}}^{(k)}$ are affected by the combined advantages of all completions in the group, while other tokens only depend on the advantage of their own completions. Under this structure, we rewrite the GRPO gradient (Eq. 2) by summing the contributions of all groups, where each group contributes a prefix term and a set of individual terms. The gradient, excluding the KL term, becomes:
$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)=\frac{1}{G}\sum_{k=1}^{K}\Bigg{[}\underbrace{\sum_{i\in\mathcal{G}_{k}}\frac{\hat{\mathcal{A}}_{i}}{|{o}_{i}|}\sum_{t\in S_{\text{pfx}}^{(k)}}g_{i,t}}_{\Delta_{\text{pfx}}^{(k)}}+\sum_{i\in\mathcal{G}_{k}}\frac{\hat{\mathcal{A}}_{i}}{|{o}_{i}|}\sum_{j\notin S_{\text{pfx}}^{(k)}}g_{i,j}\Bigg{]}$
The term $\Delta_{\text{pfx}}^{(k)}$ represents the gradient update associated with the shared prefix of group $\mathcal{G}_{k}$. If a completion does not share any prefix with others, then $\mathcal{G}_{k}$ contains only that single completion, resulting in $S_{\text{pfx}}^{(k)}=\emptyset$ and $\Delta_{\text{pfx}}^{(k)}=0$.
Since the log-probability terms $g_{i,t}$ are identical and so constant across completions in the group $S_{\text{pfx}}^{(k)}$, the gradient component $\Delta_{\text{pfx}}$ is primarily driven by the sum of normalized advantages $\sum_{i\in\mathcal{G}_{k}}\hat{\mathcal{A}}_{i}/|{o}_{i}|$. This implies that the update to a prefix mainly depends on the relative balance of advantages across all completions in group $k$. For instance, the sum can be negative, when correct completions (with positive advantages) are generally longer than incorrect ones. In such cases, the denominator $|{o}_{i}|$ for the correct completions becomes larger, reducing their contribution to the overall gradient. As a result, the model may be penalized for generating desirable and beneficial prefix tokens. This introduces a bias against shared initial tokens, such as formatting tokens or reasoning tags (<reasoning>), which are essential for the structure and correctness of completions.
Dependencies From Completion Advantage. GRPO can also induce undesirable penalization on tokens that appear after the prefix. Certain tokens, in particular formatting ones, may occur in multiple completions at varying positions and within different contexts, some correct, others incorrect, making them more susceptible to inconsistent and potentially harmful updates.
This induces potential issues in the update. For instance, if a token $\tau$ frequently appears in completions with negative advantage, its probability may be reduced, even if the token is syntactically correct and required. Additionally, further issues can arise when a completion only partially follows the expected format. For example, if a completion closes a first part correctly with a formatting token $\tau$ </reasoning> but then omits <answer>, $\tau$ may receive a low or even negative reward. This, in turn, penalizes </reasoning>, despite it being a correct token in the completion.
This occurs because the update for each token primarily depends on the advantage assigned to the entire completions in which it appears, without directly considering whether the token’s individual contribution was beneficial. We acknowledge that this phenomenon is particularly impactful in the case of formatting tokens, as illustrated in the previous examples, and in the case of shared final tokens that form a common suffix among completions.
4.2 Policy Collapse
The GRPO gradient, as discussed in Eq. (2), highlights a dependency on both the advantage and the probability scores assigned by the policy $\pi_{\theta}$. To gain a deeper understanding of how the advantage influences the learning dynamics, we analyze the average Shannon entropy of the output distribution $\pi_{\theta}$ over completions $i$ and tokens $t$ during training. The Shannon entropy is defined as $H(\mathbf{p})=-\sum_{j=1}^{n}p_{j}\ln p_{j},$ where $\mathbf{p}$ is a probability vector. For a token position $t$ in a completion $o_{i}$, we consider $\mathbf{p}=\pi_{\theta}(o_{i,t}|s_{i,t})$, and we define $\langle H\rangle_{i}$ as the average entropy across in the completion $o_{i}$.
At a generation step, let us consider a low entropy (e.g., $\langle H\rangle_{i}\ll\ln 2$)  [^2] A completion with average entropy near $\ln 2$ suggests balanced uncertainty between two choices for each output token. , indicating that the model is highly confident, i.e., it assigns almost all probability mass to a single token index $j$, with $\pi_{\theta}^{j}\approx 1$ and $\pi_{\theta}^{k}\approx\epsilon\ll 1$, where $k\neq j$. If this token leads to a negative outcome ($\hat{\mathcal{A}}_{i}<0$), GRPO penalizes it sharply, since the gradient of the GRPO loss at index $\tau$ for completion $i$ becomes:
$\nabla_{\theta}{\mathcal{J}}^{(i,\tau)}_{\text{GRPO}}(\theta)\approx\begin{cases}-\dfrac{|\hat{\mathcal{A}}_{i}|}{|o_{i}|}\cdot\nabla_{\theta}f_{\theta}^{j}\cdot(1-\epsilon)&\text{for }j\\ \;\;\dfrac{|\hat{\mathcal{A}}_{i}|}{|o_{i}|}\cdot\nabla_{\theta}f_{\theta}^{k}\cdot\epsilon&\text{for }k\neq j\end{cases}$ (3)
Note that the KL term has been omitted here, its contribution will be addressed next.
This results in a negative update for the selected token $j$, and small positive updates for all other tokens $k\neq j$, even if they are potentially syntactically or semantically incorrect. While each of these positive updates is individually small, they can accumulate over time, gradually increasing the probability of initially implausible tokens.
The effect is further amplified by GRPO’s zero-mean constraint: when most completions receive positive rewards, the few with negative rewards must carry disproportionately large negative advantages to maintain balance. These penalties can suppress correct predictions and unintentionally amplify the likelihood of undesirable alternatives, raising entropy and introducing instability. Over time, this harms output quality and increases the risk of introducing structural and semantic errors that propagate through subsequent training steps. As the model drifts away from desirable behaviors, the reward signal becomes less meaningful.
Figure 1: Training behavior of LLaMA-8B on GSM8K using two GRPO runs with $\beta=0.04$ and $\beta=10^{-6}$, both with $G=8$. The top plot shows the formatting rewards for both runs, and the bottom plots show the average entropy and KL divergence in each run.
This phenomenon is illustrated in the top plot of Figure 1, which shows the average formatting reward during the training of LLaMA-8B on GSM8K, using GRPO with $\beta=0.04$ (as originally proposed in (Shao et al. 2024)) and $\beta=10^{-6}$. At the beginning of training, the formatting reward increases more rapidly with $\beta=10^{-6}$, while it remains relatively flat with $\beta=0.04$. However, in both cases, the reward begins to drop below 9, after 580 steps for $\beta=10^{-6}$ and around step 750 for $\beta=0.04$.
KL Reacts Late to Avoid Collapse. Middle and bottom plots of Figure 1 show the average entropy $\langle H\rangle_{i}$ and average KL divergence, computed on the generated completions, for the two previous training runs of LLaMA 8B. Both KL divergence curves remain stable up to approximately step 1200, well after the degradation points, when the formatting rewards have already collapsed and entropy levels are high. Only around the KL peak, the reward curves show signs of stabilization. In contrast, the entropy curves begin to rise sharply as the formatting rewards start to decline, and continue to increase as the rewards deteriorate further.
This pattern suggests a key insight: KL divergence acts as a delayed corrective signal, rising only after collapse. In contrast, entropy tracks the policy collapse in real time, increasing as the policy loses structure. Further analyses of the entropy behaviour are provided in the appendix.
5 Method
This section introduces GTPO, which addresses the previous issues by (i) mitigating gradient conflict on shared tokens, and (ii) preventing policy collapse during training by regularizing the model behavior through the entropy.
5.1 Conflict-Aware Gradient Correction
As discussed in Section 4.1, The model can encounter gradient conflict issues, where shared tokens receive both positive and negative updates depending on the advantage assigned to each completion. To correct this issue, we propose in the following a methodology aimed at selectively masking gradient contributions for tokens involved in conflicting updates. To this end, we first identify conflict tokens, as tokens that appear in the same position, either from the beginning or the end, across completions with both positive and negative advantages. As discussed in the context of GRPO’s issues part, these tokens often receive conflicting gradient updates during training. Then, we mitigate such conflicts by correcting their gradient updates accordingly.
Conflict tokens definitions. Let $\{o_{1},\dots,o_{|G|}\}$ be a group of completions. Each completion $o_{i}$ contains a sequence of tokens $\{o_{i,1},\dots,o_{i,|o_{i}|}\}$, and is associated with an advantage value $A^{(i)}\in\mathbb{R}$. We define $G^{-}$ and $G^{+}$ as the sets of completions with $\mathcal{A}<0$ and $\mathcal{A}>0$, respectively.
Left-to-right alignment. A token $v\in\mathcal{V}$ is a forward conflict token at position $p$ if it appears at the same position in at least one completion with positive advantage and in at least one with negative advantage:
$\exists\,i\in G^{+}:\ o_{i,p}=v\quad\wedge\quad\exists\,j\in G^{-}:\ o_{j,p}=v.$
Right-to-left alignment. A token $v$ is defined as a backward conflict token at offset $r$ if it occurs at the $r$-th position from the end in at least one completion with a positive advantage and in at least one with a negative advantage:
$\exists\,i\in G^{+}:\ o_{i,|o_{i}|-r}=v\quad\wedge\quad\exists\,j\in G^{-}:\ o_{j,|o_{j}|-r}=v.$
Gradient Reweighting with conflict masks. Based on the definitions above, we construct binary masks for tokens across completions to target positions potentially affected by gradient conflicts and thus correct their updates.
We first define a forward mask $\mathcal{M}^{\mathrm{fw}}_{i}\in\{0,1\}^{|o_{i}|}$ by scanning $o_{i}$ from left to right, setting 1 over the first contiguous span of forward conflict tokens and 0 elsewhere. Analogously, the backward mask $\mathcal{M}^{\mathrm{bw}}_{i}$ is obtained by scanning from right to left, marking the first contiguous span of backward conflict tokens. A final mask $\mathcal{M}_{i}$ is then defined as: $\mathcal{M}_{i}=\mathcal{M}^{\mathrm{fw}}_{i}\lor\mathcal{M}^{\mathrm{bw}}_{i},$ highlighting only the initial and final conflict regions in each completion $o_{i}$.
After computing each $\mathcal{M}_{i}$ in the group of completions, we correct inconsistent gradient updates over the selected conflict tokens, while leaving other token updates unchanged, thought the following preliminary token-level loss:
$\mathcal{J}^{*}=\frac{1}{G}\sum_{i=1}^{G}\frac{\mathcal{A}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\lambda_{i,t}$ (4)
where $\lambda_{i,t}$ controls the update of each token based on its conflict position and the sign of the advantage $\mathcal{A}_{i}$:
$\lambda_{i,t}=\begin{cases}1&\text{if }\mathcal{M}_{i,t}=0,\\ 0&\text{if }\mathcal{M}_{i,t}=1\text{ and }\mathcal{A}_{i}<0,\\ 2&\text{if }\mathcal{M}_{i,t}=1\text{ and }\mathcal{A}_{i}>0.\end{cases}$ (5)
Intuitively, the mask disables negative gradients on conflict tokens, and instead reinforces them only if they appear in positively rewarded completions. Note that the total signal magnitude is preserved across the group: since $\sum_{i\in G^{+}}|\mathcal{A}_{i}|=\sum_{i\in G^{-}}|\mathcal{A}_{i}|$, doubling the signal for positive completions compensates for the removal of negative updates, maintaining training stability while preventing semantic drift. Additional details of the weighting scheme are provided in appendix, while its benefits are evaluated in the ablation studies (Section 6.2).
Importantly, the final mask $\mathcal{M}_{i}$ targets only the initial and final contiguous conflict tokens to preserve the semantic structure of completions. In fact, masking isolated conflict tokens in the middle could harm stability and learning, as their meaning often depends on surrounding context, and altering their gradients may be counterproductive. In contrast, the outer spans typically correspond to formatting tags, such as <reasoning> or </answer>, which are the primary source of conflict in GRPO, as discussed in Section 4.1. Focusing the correction on these regions protects structural tokens without interfering with the central part of the completion, where meaningful differences in trajectories emerge.
5.2 Entropy-Based Policy Regularization
As discussed in Section 4.2, GRPO can lead to policy collapse, where standard KL term may react too slowly. To address this, we propose entropy-based regularization terms during training. These consist of two key parts: (i) a filtering mechanism to discard unstable completions, and (ii) a regularization term that penalizes high-entropy behavior.
Completion filter. Based on the policy-collapse analysis, we observe that high-entropy completions can jeopardize training by signaling structural uncertainty, particularly in models that naturally exhibit low average entropy. Applying gradients in such cases risks amplifying uncertainty and accelerating collapse. To mitigate this, we propose filtering out high-entropy completions, focusing on models prone on collapse against this. We define $\langle H\rangle_{\text{ini}}$ as the model’s initial entropy over a set of questions, measured prior to training. If $\langle H\rangle_{\text{ini}}<\ln 2$, we assume that the model tends to produce low-entropy outputs, making it more sensitive to high-entropy completions during training. In this case, we apply an entropy-based filtering mask $\delta_{i}$ that filter out the associated advantage signal. The mask $\delta_{i}$ is formally defined as:
$\delta_{i}=\begin{cases}1,&\text{if }\langle H\rangle_{ini}>\ln 2,\\ 0,&\text{if }\langle H\rangle_{ini}<\ln 2\text{ and }\langle H\rangle_{i}>\ln 2,\\ 1,&\text{if }\langle H\rangle_{ini}<\ln 2\text{ and }\langle H\rangle_{i}\leq\ln 2.\end{cases}$ (6)
Entropy Regulatization. Inspired by PPO (Andrychowicz et al. 2021; Huang et al. 2022), we add a regularization term based on the average tokens entropy of each completion, $\langle H\rangle_{i}$, where $\gamma$ balance the importance of this term in the final loss, as shown below. Note that, based on the internal characteristics of GRPO to implicitly increase entropy over time, we decided to minimize the term. This acts as a way to reduce the model entropy over time.
The combination of these entropy-based strategies and the token-level loss defines the final GTPO objective, as:
$\mathcal{J}_{\text{GTPO}}=\frac{1}{G}\sum_{i=1}^{G}\frac{\delta_{i}\cdot\mathcal{A}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\lambda_{i,t}\;-\;\gamma\cdot\langle H\rangle_{i}$ (7)
The proposed GTPO loss does not require an additional reference model, unlike the KL divergence term in GRPO, thereby reducing the memory footprint during training. The benefits of each component, and hyperparameter of the loss are discussed and demonstrated, also through ablation studies in the following experimental section.
6 Experiments
Experimental Setup. We conducted experiments to assess the training stability and generalization performance of GTPO. All experiments were performed on LLaMA-8B (Patterson et al. 2022) and Qwen 2.5 (3B) (Yang et al. 2025a), which have trained using the training splits of GSM8K (Hendrycks et al. 2021a) and MATH (Hendrycks et al. 2021b), and evaluated after training on the corresponding test splits, and also the AIME2024 dataset (AIME 2024).
Figure 2: Training accuracy and formatting rewards (%) of GTPO and GRPO over training steps on MATH and GSM8K (top). In-distribution evaluation of models trained with GTPO, GRPO, and SFT on the corresponding test sets, using pass@k and maj@k (%) over $k$ (bottom).
Figure 3: Out-of-distribution evaluation on AIME2024, with pass@k (%) over $k$, on models trained on MATH and GSM8K.
For comparison, all models were also trained using SFT and GRPO (with both $\beta=0$ and $\beta=10^{-6}$ to assess the impact of the KL term). [^3] Additional analysis of different $\beta$ values is provided in the appendix; $\beta=10^{-6}$ was selected as it yields the best results. To further explore hyperparameter group-relative optimizations, both GRPO and GTPO were evaluated with two generation sizes: $G=8$ and $G=12$. For the entropy-based completion mask used in GTPO, we compute $\langle H\rangle_{\text{ini}}$ by evaluating the original LLM entropy on the first 100 samples of the training set.
All training was performed using a learning rate of $10^{-6}$, with the temperature set to 1.0 during the test phase. Further details of the experimental setting are in the Appendix. All experiments were conducted on 2 NVIDIA A100 GPUs.
6.1 Performance Evaluation
Training Dynamics of GTPO. In Figure 2-top (the first two rows), we compare the training stability and performance of GTPO against GRPO. Formatting and accuracy rewards are reported as percentages, where the maximum value (100%) corresponds to a reward of 10 in both. On the GSM8K dataset with LLaMA, GTPO consistently outperforms GRPO across all training steps in both accuracy and formatting metrics. On the more challenging MATH dataset, GRPO (with $G=12$ and $\beta=0$) initially achieves slightly better accuracy around the midpoint of training. However, its performance drops sharply in the second half of training due to policy collapse, affecting both accuracy and formatting. In contrast, GTPO continues to improve steadily throughout training, avoiding collapse and maintaining stable performance. For Qwen 2.5, on both GSM8K and MATH, GTPO achieves comparable or improved accuracy relative to GRPO, with only a slight decrease in formatting performance (still above 97% in all runs). Note that, consistent with the analysis in Section 4.2, GRPO training curves for Qwen 2.5 are not affected by a policy collapse, as it exhibits high-entropy behavior. Overall, GTPO demonstrates more stable and reliable training compared to GRPO across models and datasets.
In-distribution Evaluation. The trained LLMs were evaluated on the test sets of GSM8K and MATH using the pass@k (Chen et al. 2021) and maj@k (Wang et al. 2023) metrics. The former measures whether at least one of the top-$k$ completions yields a correct answer, while the latter assesses correctness via majority voting over the top-$k$ completions. Note that, to ensure a fair comparison, we evaluated GRPO on LLaMA using the model checkpoint corresponding to the training step with the highest accuracy reward, rather than the one obtained after policy collapse.
Figure 2-bottom (the last two rows) shows that GTPO consistently outperforms GRPO in almost all settings for both pass@k and maj@k, as $k$ varies from 1 to 32. This indicates that models trained with GTPO exhibit stronger self-consistency when answering questions (higher maj@k) and better coverage of the correct answer across multiple completions (higher pass@k). We also include testing comparisons with SFT, where GTPO consistently outperforms SFT in maj@k across all models and datasets, and achieves higher average performance in pass@k. Specifically, SFT surpasses GTPO only in pass@k for $k>5$ on MATH with LLaMA, but not in terms of correctness with maj@k. Interestingly, in GTPO, larger values of $G$ lead always to better performance, which is not the same for GRPO.
Figure 4: Analysis of training dynamics under different settings of entropy regularization $\gamma$: average group completion entropy (left), accuracy (center), and formatting (right). Note that $\gamma=0.1$, No $\delta_{i}$’ denotes the setting in which the entropy-based filtering term is disabled.
Figure 5: Accuracy and formatting performance of LLaMA trained on GSM8K. For GRPO, results are shown under varying KL-$\beta$ values. For GTPO, “No $\langle H\rangle_{i}$ - No $\delta_{i}$" denotes the setting in which only the Conflict-aware Gradient Correction is applied.
Out-of-distribution Evaluation. We also evaluate the trained models on a different test set (AIME2024), as shown in Figure 3, reporting pass@k scores with $k$ extended up to 64 to account for the increased difficulty of the task. For convenience, for each dataset, we select the GRPO and GTPO variants with the generation size $G$ that yielded the best performance on the in-distribution tests.
The results show that GTPO consistently outperforms both SFT and GRPO, particularly at higher values of $k$ on the MATH dataset, where the increased complexity encourages broader exploration of reasoning paths. In contrast, GRPO does not consistently benefit from this, which limits its performance gains. Interestingly, both GTPO and GRPO demonstrate stronger out-of-distribution generalization compared to SFT, suggesting a higher degree of overfitting to the in-distribution data in the latter.
6.2 Ablation Studies
Entropy-based terms. To evaluate the impact of entropy-based terms, Figure 4 shows the training curves for average completion entropy (left), accuracy (middle), and formatting (right) for LLaMA-8B on the MATH dataset. We compare different entropy regularization strengths: $\gamma=0.1$, $\gamma=0.01$, $\gamma=0.001$, $\gamma=10^{-6}$, and $\gamma=0.1$ without applying the filtering defined in Eq. 6 (denoted as “NO $\delta_{i}$”). This last setting highlights the impact of not filtering out high-entropy completions that could trigger policy collapse.
As shown in the figure, larger values of $\gamma$ lead to improved accuracy and formatting, with $\gamma=0.1$ achieving the highest accuracy overall. In contrast, when filtering is disabled ($\gamma=0.1$, No $\delta_{i}$), both accuracy and formatting collapse, highlighting the critical role of the filtering term in maintaining training stability. This behavior is further supported by the entropy plot (left side): without filtering, entropy steadily increases and remains above $\ln 2$, eventually destabilizing formatting (initially, as seen in the red region) and later affecting accuracy. In contrast, when filtering is applied, entropy remains below $\ln 2$ and gradually decreases over time.
Interestingly, the figure shows that higher values of $\gamma$ lead to both greater stabilized entropy (left plot) and improved performance. Regarding performance, in terms of accuracy and formatting, this trend can be explained by the fact that very low entropy causes the model to become overly confident, limiting its ability to explore (e.g., with $\gamma=0.001$, where entropy continues to decrease and accuracy plateaus around step 4000). In contrast, having moderate entropy promotes continued exploration during training, resulting in more diverse and informative completions.
To better understand the behaviour of the entropy curves, where larger $\gamma$ values appear to converge toward higher entropy values, when a completion receives a negative advantage, entropy tends to increase (see Eq. 4). In this context, a stronger entropy regularization term (Eq. 7) amplifies the negative gradient on the selected tokens, increasing the probability of the unselected ones. This flattens the output distribution slightly, encouraging the model to continue exploring even in the later training stages.
Conflict-Aware gradient correction. Figure 5 shows the accuracy and formatting training curves of LLaMA on GSM8K. The model is trained using GRPO with KL $\beta$ values set to 0, 0.04 (Shao et al. 2024), and $10^{-6}$, while for GTPO, we use the full version (Eq. 7) and a variant without entropy-based filtering and regularization (denoted as “No $\langle H\rangle_{i}$ - No $\delta_{i}$" in the figure). This latter configuration isolates the effect of GTPO when relying solely on the Conflict-Aware Gradient Correction component (i.e., Eq. 4). As shown in the figure, GTPO outperforms GRPO in both accuracy and formatting during the first 2,500 steps. Beyond this point, GTPO with regularization and filtering continues to maintain better performance, whereas the variant without these components begins to degrade and eventually falls below GRPO. This behavior is expected, as the absence of regularization prevents the model from balancing the impact reward signals over time. Most importantly, before the policy collapse, the use of gradient correction alone yields higher rewards than GRPO, highlighting its benefits.
7 Conclusion
In this work, we presented GTPO (Group-relative Trajectory-based Policy Optimization), a stable and effective policy optimization method for language models. GTPO addresses two key issues identified in GRPO: gradient conflicts affecting shared tokens in group of completions, and policy collapse. The core idea of GTPO is to control the divergence of completions within the same group, considering them as linked trajectories. This is achieved by mitigating gradient issues through the identification and masking of conflict tokens, and addressing policy collapse via entropy-based filtering and regularization. Through comprehensive experiments on GSM8K, MATH, and AIME2024, we demonstrated that GTPO consistently outperforms both GRPO and SFT across multiple settings.
A promising future direction will be to further investigate a theoretical minimum entropy threshold, with additional insights provided in the appendix, which may guide models toward an optimal entropy level and exploration. We believe that the insights presented in this study offer important contributions into the understanding of stable model alignment within the learning dynamics of language models, particularly in relation to entropy bounds and gradient conflicts that can arise in group-relative policy optimization paradigms.
Appendix and supplementary material for the paper :
GTPO: Trajectory-Based Policy Optimization in Large Language Models
8 Extended analysis of GRPO and GTPO
8.1 The Role of KL Divergence on GRPO
We analyze the impact of the KL divergence term on the aggregated GRPO loss by transitioning from a token-level formulation to a trajectory-level one. Recall the normalized advantage function used in GRPO:
$\hat{\mathcal{A}}_{i,t}=\hat{\mathcal{A}}_{i}=\frac{R_{i}-\bar{R}}{\operatorname{std}(R)},$ (8)
where $R_{i}$ is the scalar reward associated with trajectory $i$, $\bar{R}$ is the average reward across all $G$ trajectories in the batch, and $\operatorname{std}(R)$ denotes the standard deviation of the rewards.
Substituting Equation 8 into the GRPO objective from Equation 1, we obtain:
$\mathcal{J}_{\mathrm{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{R_{i}-\bar{R}}{\operatorname{std}(R)}-\beta\cdot\mathrm{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)$ (9)
We now turn our attention to the first term of the loss, which is defined as follows:
$\tilde{\mathcal{J}}_{\mathrm{GRPO}}(\theta):=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{R_{i}-\bar{R}}{\operatorname{std}(R)}.$ (10)
Since equation 8 is independent of $t$, it is possible to simplify $\tilde{\mathcal{J}}_{\mathrm{GRPO}}(\theta)$ as follows:
$\tilde{\mathcal{J}}_{\mathrm{GRPO}}(\theta)$ $=\frac{1}{G}\sum_{i=1}^{G}\left(\frac{R_{i}}{\operatorname{std}(R)}-\frac{\bar{R}}{\operatorname{std}(R)}\right)$
$=\frac{1}{\operatorname{std}(R)}\left(\frac{1}{G}\sum_{i=1}^{G}R_{i}-\bar{R}\right)$
$=\frac{1}{\operatorname{std}(R)}(\bar{R}-\bar{R})=0.$
Thus, the overall GRPO objective reduces to:
$\mathcal{J}_{\text{GRPO}}(\theta)=-\beta\cdot\mathrm{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right).$ (11)
From this aggregated viewpoint, the loss is entirely governed by the KL divergence term. However, it is crucial to emphasize that a zero-valued aggregated reward term does not imply that the gradient of the original GRPO loss is zero. The per-token advantages still guide parameter updates during optimization, ensuring meaningful learning even when the aggregated advantage cancels out.
8.2 GRPO Gradient
Given the GRPO objective defined in Equation 1, we aim to calculate the gradient. Similar to what we have done in GRPO Loss, we decompose the objective by focusing our attention on Eq. 10.
$\tilde{\mathcal{J}}_{\text{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$ (12)
In this case, to simplify the calculation, $\hat{\mathcal{A}}_{i}$ is kept unresolved (Eq. 8), and the term $\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$ is included to ensure the correct computation of the gradient. However, as shown in Section 3, this term is equal to 1 for every $t\in|o_{i}|$ because only a single iteration is considered, implying that $\pi_{\theta}=\pi_{\theta_{\text{old}}}$.
$\nabla_{\theta}\tilde{\mathcal{J}}_{\text{GRPO}}(\theta)$ $=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\nabla_{\theta}\left[\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\right]$
$=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{\nabla_{\theta}\left[\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right]}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$
We subsequently apply the log-derivative trick, a widely used technique in reinforcement learning, which reformulates the gradient as $\nabla_{\theta}\pi_{\theta}(x)=\pi_{\theta}(x)\nabla_{\theta}\log(\pi_{\theta}(x))$.
$\nabla_{\theta}\tilde{\mathcal{J}}_{\text{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\cdot\\ \cdot\nabla_{\theta}\left[\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t}))\right]$ (13)
Then considering that the ratio between $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ is equal to 1, we can rewrite the complete gradient as:
$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)$ $=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\nabla_{\theta}\left[\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t}))\right]$
$-\beta\cdot\nabla_{\theta}\left[\mathrm{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\right]$ (14)
How gradients affect distribution over tokens.
LLMs typically terminate with a softmax layer, which produces a probability distribution over the vocabulary. To simplify notation in what follows, we denote $(o_{i,t}|q,o_{i,<t})$ as $(t^{\prime})$, yielding the following expression for the output distribution:
$\pi_{\theta}(t^{\prime})=\frac{e^{f^{j}_{\theta}}}{\sum_{k}^{V}e^{f^{k}_{\theta}}}$ (15)
Here, $f^{j}_{\theta}$ denotes the chosen logit corresponding to position $o_{i,t}$, while $\sum_{k=1}^{V}f^{k}_{\theta}$ represents the aggregate logits over the entire vocabulary of size $V$, conditioned on the context $(q,o_{i,<t})$.
Our objective is to compute the gradient in order to analyze its influence on the probability distribution of both the selected token and the unselected alternatives.
$\nabla_{\theta}\log(\pi_{\theta}(t^{\prime}))$ $=\nabla_{\theta}\log\left(\frac{e^{f^{j}_{\theta}}}{\sum_{k}^{V}e^{f^{k}_{\theta}}}\right)$
$=\nabla_{\theta}\log\left(e^{f^{j}_{\theta}}\right)-\nabla_{\theta}\log\left(\sum_{k}^{V}e^{f^{k}_{\theta}}\right)$ (16)
The first term is just $\nabla_{\theta}e^{f^{j}_{\theta}}$; for the second, apply the chain rule:
$\nabla_{\theta}\log\left(\sum_{k}^{V}e^{f^{k}_{\theta}}\right)$ $=\frac{1}{\sum_{b}^{V}e^{f^{b}_{\theta}}}\nabla_{\theta}\left(\sum_{k}^{V}e^{f^{k}_{\theta}}\right)$
$=\sum_{k}^{V}\frac{e^{f^{k}_{\theta}}}{\sum_{b}^{V}e^{f^{b}_{\theta}}}\nabla_{\theta}f^{k}_{\theta}$
$=\sum_{k}^{V}\pi^{k}_{\theta}(t^{\prime})\nabla_{\theta}f^{k}_{\theta}$ (17)
Where $\pi^{k}_{\theta}(t^{\prime})$ represents the probability of the possible token $k$ given the context $t^{\prime}$, we can combine these as follows:
$\nabla_{\theta}\log(\pi_{\theta}(t^{\prime}))$ $=\nabla_{\theta}f^{j}_{\theta}-\sum_{k}^{V}\pi_{\theta}(k)\nabla_{\theta}f^{k}_{\theta}$
$=\nabla_{\theta}f^{k}_{\theta}(1-\pi_{\theta}(t^{\prime}))-\sum_{k\neq t}^{V}\pi_{\theta}(k^{\prime})\nabla_{\theta}f^{k}_{\theta}$ (18)
Substituting this into the GRPO gradient (ignoring the KL term for clarity) yields:
$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}$ $(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Bigg{(}\nabla_{\theta}f^{k}_{\theta}(1-\pi_{\theta}(t^{\prime}))$
$\left.-\sum_{k\neq t}^{V}\pi_{\theta}(k^{\prime})\nabla_{\theta}f^{k}_{\theta}\right)-\beta\cdot\nabla_{\theta}\left[\mathrm{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\right]$ (19)
8.3 Impact of KL $\beta$-coefficient on GRPO
Figure 10(a) presents an ablation study on the effect of the KL divergence coefficient $\beta$ during GRPO training on the MATH dataset using the Qwen model. The three plots respectively show: (left) the average entropy, (center) the accuracy rate, and (right) the formatting rate across training steps.
In the leftmost plot, we observe that with $\beta=0.04$ (blue line), the entropy steadily increases, indicating that the model becomes progressively less confident in its predictions. Conversely, smaller or null values of $\beta$ (orange for $\beta=0$ and pink for $\beta=10^{-6}$) result in a more stable or decreasing entropy, reflecting more deterministic behavior during generation.
The central plot highlights how lower values of $\beta$ lead to better accuracy. In particular, $\beta=0$ achieves the highest accuracy throughout training, while $\beta=0.04$ results in slower improvement and lower final performance. This suggests that a strong KL regularization term may overly constrain the policy and hinder learning.
Finally, the rightmost plot shows the formatting reward. Both $\beta=0$ and $\beta=10^{-6}$ achieve high formatting scores (above 98%) by the end of training, whereas $\beta=0.04$ lags behind. This confirms that a weaker KL constraint facilitates the preservation of structural consistency and formatting in model completions.
Overall, the figure demonstrates that reducing or removing the KL divergence term in GRPO ($\beta\rightarrow 0$) leads to better performance in terms of both accuracy and formatting, while avoiding the entropy inflation observed with larger $\beta$ values.
Figure 6: Average entropy (left), accuracy (center), and formatting rate (right) in QWEN trained with GRPO on MATH with different values of $\beta$ of KL
8.4 Sensitivity of GRPO to Adam Momentum
Figure 7 compares the impact of different Adam optimizer hyperparameter configurations, specifically the momentum coefficients $\alpha_{1}$ and $\alpha_{2}$, on the training dynamics of GRPO with generation size $G=8$. The two plots report the evolution of accuracy (left) and formatting score (right) over the global training steps.
The experiments are conducted using two models, Qwen and LLaMA, each evaluated under two Adam settings:
  • $\alpha_{1}=0.99999$, $\alpha_{2}=0.999999$ (used in GTPO (Dohare, Lan, and Mahmood 2023))
  • $\alpha_{1}=0.9$, $\alpha_{2}=0.95$ (the one used by original GRPO (Shao et al. 2024))
For Qwen (blue and orange curves), the optimizer settings do not significantly impact training stability: both configurations lead to consistent improvements in accuracy and formatting over time, with the GRPO originals ($\alpha_{1}=0.9$, $\alpha_{2}=0.95$) slightly outperforming in final accuracy.
In contrast, for LLaMA (pink and green curves), the choice of $\alpha_{1}/\alpha_{2}$ has a substantial effect. Using the GTPO ones ($\alpha_{1}=0.99999$, $\alpha_{2}=0.999999$) causes a sharp degradation in both accuracy and formatting after approximately 3500 steps, symptomatic of policy collapse. On the other hand, the GRPO default configuration results in more stable formatting behavior and mitigates collapse, although final formatting remains below 95%.
These results suggest that GRPO training with LLaMA is highly sensitive to Adam hyperparameters, and that careful tuning of $\alpha_{1}$ and $\alpha_{2}$ is crucial to maintain training stability and prevent collapse, particularly when using smaller values of $\beta$.
Figure 7: GRPO-GG=8, comparison between different adam $\alpha_{1}$ and $\alpha_{2}$ values.
9 Understanding the Late Response of KL Divergence in Policy Collapse Detection
To understand the different behaviors of entropy and KL divergence during training shown in Figure 1, consider the scenario in Figure 8. Assume we are training a model that initially exhibits low average entropy, specifically, $\langle H\rangle<\ln 2$. The KL divergence term is intended to keep the new policy (the model under training) close to the reference model by penalizing deviations.
Consider a prompt $q$,where both models generate the same first two tokens: $t_{0}$ and $t_{1}$. This implies that the conditional distributions agree at that step: $\pi_{\text{new}}(t_{1}\mid t_{0})\approx\pi_{\text{ref}}(t_{1}\mid t_{0}),$ and the KL divergence is near zero.
Next, the new model selects a third token $t_{2}$, while the reference model would have instead generated a different token $t_{2}^{\prime}$, causing the KL divergence to increase at this step.
At this point, the new model’s entropy is high. Repeated negative updates during training may have inadvertently boosted the probability of weak or incoherent tokens like $t_{2}$, flattening the output distribution and reducing confidence. In contrast, the reference model remains sharper, confidently selecting $t_{2}^{\prime}$.
Figure 8: Divergence point where the new policy deviates from the reference.
(a) Creation of Conflict Tokens.
(b) Multiplication between Forward and Backward Mask.
Figure 9: Trajectory-aware Conceptualization
Because $t_{2}$ is weakly conditioned on the prior context, the next token $t_{3}$ is sampled from an even flatter distribution. The model, having lost semantic alignment, now samples among many equally probable tokens, most of which are uninformative or incoherent. This creates a compounding effect: once the trajectory becomes unstable, each uncertain decision amplifies uncertainty in the following steps, leading to complete structural collapse. Now, consider feeding the full sequence $q,t_{0},t_{1},t_{2}$ to the reference model. Since $t_{2}$ is not a token it would naturally produced, the reference model finds itself in a semantically incoherent context. As a result, its output distribution also becomes high-entropy. In this state, both models assign similarly flat distributions to the next token. This leads to a key consequence: although the new model diverged at $t_{2}$, the KL divergence does not increase significantly in the following steps. Since KL is computed token by token, and both models are now equally uncertain, their distributions appear similar, despite the fact that the new model has collapsed. Entropy, by contrast, directly reflects the model’s internal uncertainty. As coherence deteriorates, entropy rises, signaling the model’s loss of confidence and structure. Unlike KL, entropy does not depend on comparison with a reference model: it directly reveals how dispersed the model’s beliefs are.In summary, entropy provides an early signal of policy collapse by rising with incoherent outputs, while KL divergence reacts later, only once the model confidently diverges from the reference.
10 Trajectory-aware Conceptualization
Figure 9(a) illustrates the procedure for identifying conflict tokens within a group of completions generated in response to a prompt. In this example, the prompt “What is 2 + 2?” produces several completions, each composed of a sequence of tokens and associated with either a positive (✓) or negative (✗) scalar advantage, reflecting the correctness and formatting quality of the generated output.
Tokens that appear exclusively in completions associated with advantages of the same sign, either all positive or all negative, and do not occupy the same position across completions with opposite advantages, are defined as non-conflict tokens.
As illustrated in the orange box, the numerical tokens “4” and “3” are each present in completions with different reward signs (positive and negative, respectively), but they appear in different completions and at distinct positions. Since they are not shared between completions with opposing advantages at the same aligned position (either from the start or the end), they are not considered conflict tokens and their gradients are unaffected by conflicting signals.
In contrast, conflict tokens, highlighted in the blue box, are those that appear in the same position across completions with both positive and negative advantages. These positions are evaluated either from the beginning (forward alignment) or from the end (backward alignment) of the completions. Typical examples include formatting tokens such as <reasoning>, lexical fragments like “ing”, and closing markers like </answer>. Due to their presence across completions with conflicting rewards, these tokens are subject to contradictory gradient updates during training, which can hinder convergence or disrupt the structural integrity of generated outputs.
To mitigate this, GTPO identifies such tokens using position-based alignment and applies a conflict-aware masking strategy. Specifically, it suppresses negative gradient contributions and amplifies positive ones, ensuring that structurally important or frequently reused tokens are not inadvertently penalized. This approach maintains consistency in output formatting and improves the stability and effectiveness of policy optimization during training.
Figure 9(b) illustrates how GTPO constructs and applies forward and backward conflict masks to identify and handle conflict tokens differently depending on the sign of the completion’s advantage.
On the left, we observe a completion with positive advantage (green background). The forward mask highlights the initial span of conflict tokens, namely the ones that compose <reasoning>, with binary value 1, and sets the remaining positions to 0. Similarly, the backward mask is applied from the end, marking </answer> as part of the final span of backward conflict tokens. The final mask is computed as the element-wise logical OR between the forward and backward masks, thereby identifying the full set of conflict positions. In the case of positive advantage, GTPO assigns a weight of $\lambda=2$ to these tokens, amplifying their positive update signal.
On the right, the same masking procedure is applied to a completion with negative advantage (red background). The same tokens appear as conflict tokens at the beginning and the end of the sequence, resulting in identical forward and backward masks. However, in this case, GTPO sets the update weight to $\lambda=0$ for all masked positions, effectively suppressing the negative gradient updates on those tokens. The remaining (non-conflict) tokens are assigned a default weight $\lambda=1$ in both cases.
This mechanism ensures that formatting tokens commonly shared across completions, such as <reasoning> and </answer>, are not penalized by negative completions while being reinforced in positively rewarded ones, preserving the structural integrity of the output.
11 Group-relative Trajectory-aware Policy Optimization: Objective and Gradient
In this section, we formalize the Conflict-Aware Gradient Correction of GTPO introduced in the main paper, provide the full expression of the token-level loss function and include the gradient.
GTPO builds upon GRPO but introduces a fine-grained, token-level control over gradient updates. Specifically, each token within a completion is weighted by a coefficient $\lambda_{i,t}$, which depends on its position relative to conflict regions (as defined by the forward and backward masks) and on the sign of the associated advantage $A_{i}$. Additionally, an entropy-based filtering term $\delta_{i}$ is used to discard unstable completions in models prone to policy collapse.
The resulting loss function can be expressed as:
$\mathcal{J}^{*}$ $:=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t\in o_{i}}\lambda_{i,t}A_{i}$
$=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\left(\sum_{t\in c_{i}}\lambda_{i,t}A_{i}+\sum_{t\in u_{i}}\lambda_{i,t}A_{i}\right)$
Where $o_{i}$ denotes the set of tokens in each completion, $c_{i}$ represents the subset of conflicting tokens, and $u_{i}$ denotes the subset of non-conflicting tokens within each completion. The notation $|\cdot|$ indicates the cardinality (number of elements) of each set and $\lambda_{i,t}$ is a parameter designed to account for conflict and non-conflict formulation.
We would like to retain the characteristics of GRPO for non-conflict tokens, so we define lambda as:
$\lambda_{i,t}=\begin{cases}\lambda_{i},&\text{if t}\in c_{i}\\[4.0pt] 1,&\text{if t}\in u_{i}\end{cases}$
$\mathcal{J}^{*}$ $=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\left(|c_{i}|\lambda_{i}A_{i}+\left(|o_{i}|-|c_{i}|\right)A_{i}\right)$
$=\frac{1}{G}\left(\sum_{i=1}^{G}\frac{|c_{i}|}{|o_{i}|}\lambda_{i}A_{i}+\sum_{i=1}^{G}A_{i}-\sum_{i=1}^{G}\frac{|c_{i}|}{|o_{i}|}A_{i}\right)$
$=\frac{1}{G}\left(\sum_{i=1}^{G}\frac{|c_{i}|}{|o_{i}|}A_{i}(\lambda_{i}-1)+\sum_{i=1}^{G}A_{i}\right)$
First of all, given the GRPO’s zero-mean constraint, $\sum_{i}^{G}A_{i}=0$, we split the sum into two parts: one considering only the positive completions ($G^{+}$), and the other considering only the negative ones ($G^{-}$).
$\mathcal{J}^{*}$ $=\frac{1}{G}\left(\sum_{i=1}^{G^{+}}\frac{|c_{i}|}{|o_{i}|}A_{i}(\lambda_{i}-1)+\sum_{i=1}^{G^{-}}\frac{|c_{i}|}{|o_{i}|}A_{i}(\lambda_{i}-1)\right)$
To prevent inadvertently penalizing tokens that contribute positively in other contexts, we refrain from assigning negative rewards to tokens involved in conflicting cases. These tokens frequently appear in completions that receive positive ratings and are not intrinsically responsible for the negative assessments.
$\lambda_{i}=\begin{cases}\lambda,&\text{if i}\in G^{+}\\ 0&\text{if i}\in G^{-}\end{cases}$
$\mathcal{J}^{*}$ $=\frac{1}{G}\left(\sum_{i=1}^{G^{+}}\frac{|c_{i}|}{|o_{i}|}A_{i}(\lambda-1)-\sum_{i=1}^{G^{-}}\frac{|c_{i}|}{|o_{i}|}A_{i}\right)$
Now, let us factor out the sign of each advantage and consider its absolute value.
$\mathcal{J}^{*}$ $=\frac{1}{G}\left(\sum_{i=1}^{G^{+}}\frac{|c_{i}|}{|o_{i}|}|A_{i}|(\lambda-1)+\sum_{i=1}^{G^{-}}\frac{|c_{i}|}{|o_{i}|}|A_{i}|\right)$
It is evident that by setting $\lambda=2$, we ensure equal contribution from each completion, regardless of whether it is positive or negative.
The resulting equation for the aggregated loss is:
$\mathcal{J}^{*}$ $=\frac{1}{G}\sum_{i=1}^{G}\frac{|c_{i}|}{|o_{i}|}|A_{i}|\$(20)
Instead the resulting equation for the token level loss is:
$\mathcal{J}^{*}=$ $\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=0}^{o_{i}}\lambda_{i,t}A_{i}$ (21)
where:
[…]
Gradient:
To calculate the gradient of our proposed loss, we should start from equation 21, then we will apply the same tricks used to calculate GRPO loss (eq. 12).
$\nabla_{\theta}\mathcal{J}^{*}=$ $\;\frac{1}{G}\sum_{i=1}^{G^{+}}\frac{A_{i}}{|o_{i}|}\left(2\sum_{t\in c_{i}}g_{i,t}+\sum_{t\in u_{i}}g_{i,t}\right)$
$+\frac{1}{G}\sum_{i=1}^{G^{-}}\frac{A_{i}}{|o_{i}|}\sum_{t\in u_{i}}g_{i,t}$ (23)
where:
$g_{i,t}=\nabla_{\theta}\left[\log\left(\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\right)\right]$ (24)
12 Ablation of $\langle H\rangle_{\text{i}}$ Threshold in GTPO
Figure 10(a) illustrates the effect of different thresholds for the average entropy $\langle H\rangle_{\text{i}}$ in GTPO training, using the LLaMA model on the GSM8K dataset. The three plots report the evolution of entropy (left), accuracy (center), and formatting rate (right) over training steps.
Each curve corresponds to a different entropy filter configuration:
  • Orange line: strict filtering with $\langle H\rangle_{\text{i}}<0.7$ (2 alternative tokens with the same probability)
  • Blue line: moderate filtering with $\langle H\rangle_{\text{i}}<1.05$ (3 alternative tokens with the same probability)
  • Green line: loose filtering with $\langle H\rangle_{\text{i}}<2.00$ (4 alternative tokens with the same probability)
In the entropy plot (left), we observe that stricter filtering (orange and blue) leads to a consistent reduction in average entropy during training, whereas the looser configuration (green) exhibits a growing trend, approaching the instability threshold of $\ln 2$. This suggests that high-entropy completions left unfiltered can dominate the learning signal and destabilize the policy.
The accuracy plot (center) confirms this trend: models trained with stronger entropy filtering (<0.7<0.7 and <1.05<1.05) achieve substantially higher final accuracy, with the strictest setting yielding the best performance. In contrast, training with $\langle H\rangle_{\text{ini}}<2.0$ leads to stagnation and ultimately lower accuracy.
The formatting rate plot (right) further highlights the benefits of filtering. Without entropy-based filtering (green line), formatting rapidly degrades, suggesting collapse, while stricter filtering helps maintain structural consistency throughout training.
Overall, the figure demonstrates that entropy-aware filtering is essential for stabilizing GTPO, particularly when using models such as LLaMA that are sensitive to noisy or uncertain completions. Proper calibration of $\langle H\rangle_{\text{i}}$ is thus critical to avoiding policy collapse and ensuring both performance and formatting integrity.
(a) Average entropy (left), accuracy (center), and formatting rate (right) in LLaMA trained with GTPO on GSM8K with different values of $\langle H\rangle_{i}$.
(b) Average entropy (left), accuracy (center), and formatting rate (right) in GRPO, GTPO and GTPO applying different $\langle H\rangle_{i}$ thresholds only to completions with negative advantages.
(c) Average entropy (left), accuracy (center), and formatting rate (right) for LLaMA (GSM8K), in GRPO, GRPO with $\langle H\rangle_{i}<1.05$ and $\delta_{i}=0.1$ and GTPO with $\langle H\rangle_{i}<1.05$ and $\delta_{i}=0.1$.
(d) Impact of GTPO group size ($G=8$ vs. $G=12$) on GSM8K (top) and MATH (bottom) for QWEN and LLaMA. Metrics shown over training steps: accuracy, formatting reward, entropy, and mean conflict (left to right).
Figure 10: Ablation results comparing entropy, accuracy, formatting rate, and conflict under different strategies and hyperparameters for GTPO and GRPO across GSM8K and MATH.
13 Entropy-Based Filtering: Global vs. Negative-Only
Figure 10(b) presents a comparative analysis of training dynamics when entropy-based filtering is applied exclusively to completions with negative advantages. The figure reports average entropy (left), accuracy (center), and formatting rate (right) for GRPO and multiple variants of GTPO under different entropy threshold conditions.
The baseline GRPO curves (blue for $\beta=10^{-6}$, orange for $\beta=0$) exhibit moderate to high entropy, with degraded formatting and suboptimal accuracy, particularly in the $\beta=10^{-6}$ setting. In contrast, GTPO variants introduce entropy thresholds, applied solely to negatively rewarded completions, to selectively suppress gradient updates from high-entropy, unreliable trajectories that may contribute to policy collapse.
Three GTPO-OnlyNeg configurations are shown:
  • Green: relaxed threshold $\langle H\rangle_{i}<2.08$ (4 alternative tokens with the same probability)
  • Yellow: intermediate threshold $\langle H\rangle_{i}<1.10$ (3 alternative tokens with the same probability)
  • Blue: strict threshold $\langle H\rangle_{i}<0.7$ (2 alternative tokens with the same probability)
Additionally, the pink curve represents the standard GTPO policy, where filtering is applied across all completions (positive and negative) under the strict threshold $\langle H\rangle_{i}<0.7$.
The entropy plot (left) shows that restricting filtering to only negative completions still significantly reduces overall entropy, especially with tighter thresholds. The accuracy plot (center) indicates that all GTPO-OnlyNeg configurations outperform GRPO, with the strictest setting (<0.7<0.7) achieving the highest final accuracy. Notably, the pink curve (GTPO with global filtering) slightly outperforms compared to its selective counterpart, because this setting lowers the entropy but the reached minimum entropy plateau is higher than the OnlyNegative ones. This allows higher exploration and so better accuracy at higher training step.
The formatting rate plot (right), shows that applying entropy filtering only to negative completions stabilize structural consistency throughout training (and also improve GTPO), even if also all the other configurations stays above 97%.
Overall, this experiment demonstrates that selectively applying entropy filtering to negative completions (GTPO-OnlyNeg) improves formatting stability and mitigates policy collapse more aggressively, achieving the highest formatting scores. However, the standard GTPO configuration, where entropy filtering is also applied to positively rewarded completions, maintains a higher entropy floor, enabling broader exploration. This leads to superior final accuracy while preserving high formatting consistency (above 97%), striking a better balance between structural integrity and generalization.
14 Effect of Regularization and Filtering on GRPO
Figure 10(c) evaluates the combined impact of entropy-based filtering and entropy regularization across different training strategies. Specifically, it compares three configurations for LLaMA (GSM8K): (i) standard GRPO without entropy control ($\beta=0$), (ii) GRPO enhanced with both filtering ($\langle H\rangle_{i}<1.05$) and entropy regularization ($\gamma=0.1$), and (iii) GTPO with the same entropy constraints.
The leftmost plot shows the average token-level entropy. GRPO without control collapses rapidly, confirming its vulnerability to policy collapse. Incorporating entropy-based filtering and regularization in GRPO slows down this collapse, maintaining moderate entropy throughout training. However, GTPO demonstrates the most robust behavior, keeping entropy stable and above the collapse threshold.
In the center plot, we observe accuracy trends. GRPO without entropy control shows limited improvement, whereas adding entropy filtering and regularization significantly enhances performance. Notably, GTPO outperforms both configurations, highlighting the synergy between conflict-aware masking and entropy-based stabilization.
Finally, the rightmost plot presents formatting accuracy. GRPO without entropy control exhibits a severe collapse in structural formatting. Entropy filtering and regularization partially restore structure in GRPO, but GTPO achieves the highest formatting consistency, preserving nearly perfect output formatting.
Overall, this figure demonstrates that combining filtering ($\delta_{i}$) and entropy regularization ($\gamma$) enhance the performances of GRPO, but is most effective when paired with GTPO’s gradient reweighting mechanism.
15 Effect of GTPO Group Size on Training Dynamics
Figure 10(d) compares GTPO with two group sizes (G=8G{=}8 vs. G=12G{=}12) on GSM8K (top row) and MATH (bottom row) for QWEN and LLaMA. Each row reports, left to right: accuracy, formatting, entropy, and mean conflict.
GSM8K (top).
  • Accuracy. Larger groups help both models: QWEN–12G attains the best final accuracy, followed by LLaMA–12G. Both 8G settings converge more slowly and to lower plateaus.
  • Formatting. All runs exceed  $98\%$. Larger groups stabilize formatting earlier, with QWEN–12G reaching the highest levels.
  • Entropy. LLaMA maintains lower entropy than QWEN. Increasing the group size raises entropy slightly for LLaMA but not for Qwen.
  • Mean conflict. Mean token-level conflict increases with the larger group (12G >> 8G). QWEN shows higher conflict than LLaMA, consistent with its higher entropy and greater trajectory diversity. Under GTPO, this added diversity is beneficial and does not destabilize formatting.
MATH (bottom).
  • Accuracy. The advantage of larger groups persists: LLaMA–12G achieves the top accuracy, with LLaMA–8G close behind; QWEN benefits from 12G but remains lower than LLaMA on this benchmark.
  • Formatting. QWEN consistently attains the highest formatting scores, especially with 12G (near-perfect early in training). LLaMA improves with 12G and requires more steps to match high formatting.
  • Entropy. As on GSM8K, LLaMA runs remain markedly lower-entropy than QWEN.
  • Mean conflict. Conflict grows with group size; the effect is strongest for QWEN–12G, reflecting increased trajectory diversity. LLaMA shows more moderate conflict levels, with limited sensitivity to group size.
In summary, larger groups (G=12G{=}12) improve accuracy and accelerate formatting stabilization across datasets and models.
16 Training Details and Code
16.1 Reward Settings
The reward used in this paper is composed of two components: the first one refers to the format of the model’s answer, and the second one refers to the accuracy of the model’s answer. The first component assesses formatting: the model receives a score of 10 if it includes all predefined special formatting tokens specified in the prompt (see box 16.1). If only a subset of the reasoning or answer tokens are present, the model is awarded 1 point; otherwise, it scores 0. The second component evaluates accuracy: the model obtains a score of 10 if the correct target answer is correctly enclosed within the answer tokens; otherwise, it receives 0. The final reward is the sum of these two components and is provided to guide the learning algorithms.
16.1 PROMPT You are a helpful assistant for solving math problems. Given a question, first think step by step between <reasoning> and </reasoning>. Then, give the final answer between <answer> and </answer>.
16.2 Training Settings
Listing 1 reports the main training parameters used for GRPO, GTPO and SFT, for both LLaMA and Qwen models. To implement GTPO we used Unsloth (Daniel Han and team 2023).
Implementation Details of Conflict Masks.
Function build_conflict_mask (Listing 2).
This routine constructs the conflict-aware token weights used to realize the forward/backward masking described in Section 4.1. Given token ids $\texttt{ids}\in\mathbb{N}^{G\times L}$, a validity mask $\texttt{mask}\in\{0,1\}^{G\times L}$, and Boolean flags marking, per completion, positive/negative advantages ($\texttt{pos\_flags},\texttt{neg\_flags}\in\{0,1\}^{G}$), it returns: (i) a final per-token weight mask $\texttt{final\_mask}\in\mathbb{R}^{G\times L}$ (implementing $\{\lambda_{i,t}\}$); (ii) a binary conflict indicator $\texttt{total\_conflict}\in\{0,1\}^{G\times L}$; and (iii) the number of conflict positions per completion $\texttt{n\_conflict\_seq}\in\mathbb{N}^{G}$.
Active subset. Completions with zero advantage are ignored: the function selects the active index set $\mathcal{G}_{\mathrm{act}}=\{i:\texttt{pos\_flags}[i]\vee\texttt{neg\_flags}[i]\}$. If $|\mathcal{G}_{\mathrm{act}}|=0$, it returns the identity mask (mask.float()), zero conflicts, and zero counts.
Forward (left-to-right) conflict detection. For the active completions, the code encodes each token-position pair into a unique key
$k(i,t)=t\cdot V+\texttt{ids}[i,t],\qquad V=\texttt{vocab\_size},$
and builds presence maps for positive and negative groups via bincount, clamped to $\{0,1\}$. A forward conflict at step $t$ is flagged whenever the same $(\text{token},t)$ appears in at least one positive and at least one negative completion:
$\texttt{conflict\_raw}(t)=\;$ $\mathbf{1}\left\{\exists i\in G^{+}:\,o_{i,t}=v\right\}$
 $\wedge\;\mathbf{1}\left\{\exists j\in G^{-}:\,o_{j,t}=v\right\}$
To restrict to the first contiguous span of conflicts from the start, it applies $\texttt{initial\_run}=\mathrm{cumprod}(\texttt{conflict\_raw})$, which zeroes any conflict after the first 0. The forward conflict mask is $\texttt{conflict\_fwd}=\texttt{initial\_run}\wedge\texttt{valid\_mask}$.
Forward weights. The per-token forward weight mask_fwd encodes the reweighting rule of Eq. (5): for conflict positions, positive-advantage completions receive a multiplicative factor W_RAW (amplification), while negative-advantage completions are nulled; non-conflict positions get weight 1. Formally, for completion $i$, token $t$:
[…]
and it is multiplied by mask_use to zero out padded positions.
Backward (right-to-left) conflict detection. Suffix alignment is handled by reversing sequences and correcting for per-sequence padding. Let $L$ be the max length, seq_len the valid length per completion, and $\texttt{pad\_len}=L-\texttt{seq\_len}$. The code flips ids and rotates indices to align the last valid token to the right:
$\texttt{ids\_rev}[i,t]\;=\;\text{FlipShift}(\texttt{ids}[i,\cdot],\ \texttt{pad\_len}[i]).$
It then repeats the same keying-and-bincount procedure on ids_rev, producing conflict_raw_rev and its first contiguous run initial_run_rev (backward conflicts). The corresponding weight mask mask_bw_rev is built with the same positive/negative rule as above, then unrotated and re-flipped back to the original order to obtain mask_bw_use, again masked by validity.
Final mask and diagnostics. The final per-token weight is the elementwise product
$\texttt{final\_mask}\;=\;\texttt{mask\_fwd}\ \odot\ \texttt{mask\_bw\_use},$
so that only tokens inside the initial forward/backward conflict spans are reweighted (amplified/suppressed) while all others retain weight 1. The binary conflict indicator is $\texttt{total\_conflict}=\texttt{conflict\_fwd}\ \vee\ \texttt{conflict\_rev}$, and the per-completion count n_conflict_seq is the sum over valid positions, clamped to at least 1 to avoid division-by-zero in subsequent normalizations.
Return to full layout. All masks and counts computed on the active subset are scattered back into the original $G\times L$ layout using active_idx; inactive rows receive the identity mask, zero conflicts, and zero counts. The function returns:
$(\ \texttt{final\_mask},\ \texttt{total\_conflict},\ \texttt{n\_conflict\_seq}\ ).$
Implementation Details of the GTPO Loss
The code in the listing 3 implements the GTPO objective at token level and a custom autograd.Function that computes gradients efficiently over grouped completions. It realizes the conflict-aware reweighting $\lambda_{i,t}$ and the entropy-based policy regularization ($\delta_{i}$ gating and entropy penalty) described in the main text.
Token-level loss gtpo_compute_loss.
Given logits of the current policy (new_logits) and reference logits for monitoring (old_logits), input token ids input_ids, a validity mask mask, the per-token conflict weight vector $\delta\equiv\{\lambda_{i,t}\}_{t}$ (named delta in code), and a scalar completion advantage $A_{i}$ (named advantages), the function computes
$\text{loss}_{\text{token}}(t)\;=\;-\big{(}\underbrace{\tfrac{\pi_{\theta}(y_{t}\!\mid\!s_{t})}{\pi_{\theta}(y_{t}\!\mid\!s_{t})}}_{\textstyle\text{ratio}=1\;\text{(carries grad)}}\big{)}\,A_{i}\,\lambda_{i,t},$
and averages it over valid tokens with mask. Concretely:
  1. 1.
    It forms the token log-probabilities $\log p_{\theta}(y_{t}\!\mid\!s_{t})$ via a gather over new_logits and a $\log\text{sumexp}$.
  2. 2.
    It sets ratio $\equiv 1$ as $\exp(\log p-\text{stopgrad}(\log p))$, so the forward value is 1 while preserving the policy gradient signal.
  3. 3.
    It builds the token loss $(\texttt{ratio}\cdot A_{i}\cdot\lambda_{i,t})$ and averages over valid tokens (mask).
  4. 4.
    It computes a tokenwise KL proxy between old and new distributions without gradient (for logging/monitoring), and returns the mean KL.
  5. 5.
    It normalizes the per-sequence loss by the number of conflict positions n_conflict (clamped to 1) to preserve signal magnitude when masking is active, then averages across sequences.
The arguments beta, rewards, and std_reward are kept for API symmetry/monitoring; KL is not added to the loss (consistent with GTPO).
Custom autograd UnslothEfficientGTPO.
The forward receives hidden states for the current and old policies with shape $B,G,L_{\text{full}},H]$ (batch, group size, sequence length, hidden size), the lm_head (projection to vocabulary), completion ids/masks, advantages, and optimizer scalars. It performs:
  1. 1.
    Per-completion entropy & KL. For each completion $(b,g)$, it projects hidden states to logits and calls a helper (not shown) to compute average entropy $\langle H\rangle_{i}$ (and KL for monitoring).
  2. 2.
    Entropy-based gating and penalty. It realizes the completion filter $\delta_{i}$ and the entropy penalty by modifying the advantages:
    $A_{i}\;\leftarrow\;\begin{cases}A_{i}-\eta\,\langle H\rangle_{i},&\text{if }\langle H\rangle_{i}\leq\texttt{ENT\_THRESHOLD},\\ 0,&\text{otherwise},\end{cases}$
    where $\eta=\texttt{ENT\_SCALE}$. Thus, high-entropy completions are filtered out ($\delta_{i}=0$), while low-entropy ones receive a small entropy penalty added directly to the advantage, matching the formulation of the main objective with $\delta_{i}$ and $\gamma$.
  3. 3.
    Conflict-mask construction. It builds the forward/backward conflict masks from token alignments across the group via build_conflict_mask, producing:
[…]
    • a diagnostic conflict_mask (positions identified as conflict),
    • a count n_conflict per completion (used for normalization).
  4. 4.
    Gradient w.r.t. new hidden states. For each completion, it computes logits from hidden states and invokes gtpo_compute_loss; then it uses torch.func.grad_and_value (wrapped in torch.compile) to obtain both the gradient (only w.r.t. new hidden states) and the auxiliary scalars (loss, token count, mean KL). These gradients are accumulated into a tensor $\partial\mathcal{L}/\partial\texttt{\_new\_hidden\_states}$ saved in the context.
The backward simply returns the saved gradient for _new_hidden_states and None for all other inputs, ensuring that only the current policy is updated.
Mapping to the objective.
The implemented loss corresponds to
$\mathcal{J}_{\text{GTPO}}\;=\;\frac{1}{G}\sum_{i=1}^{G}\frac{\delta_{i}A_{i}}{|\mathcal{O}_{i}|}\sum_{t\in\mathcal{O}_{i}}\lambda_{i,t}\;-\;\gamma\,\langle H\rangle_{i},$
with the following realizations in code:
  • $\lambda_{i,t}$ $\rightarrow$ delta (a.k.a. conflict mask) inside gtpo_compute_loss;
  • $\delta_{i}$ (filter) $\rightarrow$ zeroing the advantage when $\langle H\rangle_{i}>$ ENT_THRESHOLD;
  • $\gamma\,\langle H\rangle_{i} \rightarrow$ subtracting $\texttt{ENT\_SCALE}\cdot\langle H\rangle_{i}$ from $A_{i}$ when not filtered;
  • KL is computed for monitoring only and not added to the objective.
Normalizations by the number of valid tokens and by the conflict count ensure that the magnitude of the learning signal is preserved when conflict masking disables negative updates and doubles positive ones at conflict positions.
1model_name: "meta-llama/meta-Llama-3.1-8B-Instruct" / "Qwen/Qwen2.5-3B-Instruct"
2max_seq_length: 5500
3max_prompt_length: 4000
4lora_rank: 128 / 64 (for Qwen)
5load_in_4bit: true
6gpu_memory_utilization: 0.4
7target_modules:
8 - "q_proj"
9 - "k_proj"
10 - "v_proj"
11 - "o_proj"
12 - "gate_proj"
13 - "up_proj"
14 - "down_proj"
15random_seed: 3407
16warmup_ratio: 0.005
17learning_rate: 1e-6
18adam_beta1: 0.999999 / 0.9 (GRPO and SFT)
19adam_beta2: 0.999999 / 0.95 (GRPO and SFT)
20weight_decay: 0.1
21beta: 0.0 - 0.04 - 10e-6 (for GRPO) / None (for GTPO and SFT)
22lr_scheduler_type: "cosine"
23optimizer: "paged_adamw_8bit"
24logging_steps: 1
25per_device_train_batch_size: 1
26gradient_accumulation_steps: 1
27num_generations: "8/12" / "1" (for SFT)
28num_train_epochs: 1
29num_iterations: 1
30save_steps: 500
31max_grad_norm: 0.1
32report_to: ["wandb"]
Listing 1: Training settings for GTPO, GRPO and SFT.
1def build_conflict_mask(
2 ids: torch.Tensor, # (G, L)
3 mask: torch.Tensor, # (G, L)
4 pos_flags: torch.Tensor, # (G,) bool (advantage > 0)
5 neg_flags: torch.Tensor, # (G,) bool (advantage < 0)
6 vocab_size: int,
7):
8 """
9 Builds ‘forward_mask‘, ‘backward_mask‘, and ‘final_mask‘ masks,
10 ignoring completions with zero advantage.
11 Steps:
12 1. Filter active completions.
13 2. Compute forward and backward conflicts.
14 3. Reconstruct original batch layout.
15 """
16
17 G, L = ids.shape
18 device = ids.device
19
20 active_flags = pos_flags | neg_flags
21 active_idx = active_flags.nonzero(as_tuple=True)[0]
22 G_act = active_idx.numel()
23
24 if G_act == 0:
25 mask_default = mask.float()
26 conflict_zero = torch.zeros_like(mask, dtype=torch.float)
27 seq_count_zero = torch.zeros(G, dtype=torch.long, device=device)
28 return mask_default, conflict_zero, seq_count_zero
29
30 ids_use = ids[active_idx]
31 mask_use = mask[active_idx]
32 pos_use = pos_flags[active_idx]
33 neg_use = neg_flags[active_idx]
34
35 step_idx = torch.arange(L, device=device).unsqueeze(0)
36 keys = (step_idx * vocab_size + ids_use).view(-1)
37
38 rpt_pos = pos_use.repeat_interleave(L)
39 rpt_neg = neg_use.repeat_interleave(L)
40
41 freq_pos = torch.bincount(keys[rpt_pos], minlength=vocab_size * L).clamp_max(1)
42 freq_neg = torch.bincount(keys[rpt_neg], minlength=vocab_size * L).clamp_max(1)
43
44 conflict_raw = (freq_pos.bool() & freq_neg.bool())[keys].view(G_act, L)
45
46 valid_mask = (ids_use != PAD_ID)
47 initial_run = torch.cumprod(conflict_raw.to(torch.int), dim=1).bool()
48 conflict_fwd = initial_run & valid_mask
49
50 w_conf_fwd = torch.where(conflict_fwd,
51 torch.tensor(W_RAW, device=device),
52 torch.ones((), device=device))
53 mask_fwd = torch.where(
54 ~conflict_fwd,
55 1.0,
56 torch.where(pos_use.unsqueeze(1), w_conf_fwd, 0.0),
57 ) * mask_use
58
59 seq_len = mask_use.sum(-1)
60 pad_len = L - seq_len
61
62 ids_flip = ids_use.flip(-1)
63 idx_shift = (step_idx + pad_len.unsqueeze(1)) % L
64 ids_rev = ids_flip.gather(1, idx_shift)
65
66 step_rev = torch.arange(L, device=device).flip(0).unsqueeze(0)
67 keys_rev = (step_rev * vocab_size + ids_rev).view(-1)
68
69 freq_pos_r = torch.bincount(keys_rev[rpt_pos], minlength=vocab_size * L).clamp_max(1)
70 freq_neg_r = torch.bincount(keys_rev[rpt_neg], minlength=vocab_size * L).clamp_max(1)
71
72 conflict_raw_rev = (freq_pos_r.bool() & freq_neg_r.bool())[keys_rev].view(G_act, L)
73
74 initial_run_rev = torch.cumprod(conflict_raw_rev.to(torch.int), dim=1).bool()
75 conflict_rev = initial_run_rev & valid_mask
76
77 w_conf_rev = torch.where(conflict_rev,
78 torch.tensor(W_RAW, device=device),
79 torch.ones((), device=device))
80 mask_bw_rev = torch.where(
81 ~conflict_rev,
82 1.0,
83 torch.where(pos_use.unsqueeze(1), w_conf_rev, 0.0),
84 )
85
86 idx_unshift = (step_idx - pad_len.unsqueeze(1)) % L
87 mask_bw_rot = mask_bw_rev.gather(1, idx_unshift)
88 mask_bw_use = mask_bw_rot.flip(-1) * mask_use
89
90 final_mask_use = mask_fwd * mask_bw_use
91 total_conflict_use = (conflict_fwd | conflict_rev) & mask_use.bool()
92 n_conflict_seq_use = total_conflict_use.sum(-1).long().clamp_min(1)
93
94 final_mask = mask.float().clone()
95 total_conflict = torch.zeros_like(mask, dtype=torch.float)
96 n_conflict_seq = torch.zeros(G, dtype=torch.long, device=device)
97
98 final_mask[active_idx] = final_mask_use
99 total_conflict[active_idx] = total_conflict_use.float()
100 n_conflict_seq[active_idx] = n_conflict_seq_use
101
102 return final_mask, total_conflict, n_conflict_seq
Listing 2: Function build_conflict_mask
1def gtpo_compute_loss(
2 old_logits: torch.Tensor,
3 new_logits: torch.Tensor,
4 input_ids: torch.Tensor,
5 mask: torch.Tensor,
6 delta: torch.Tensor,
7 beta: float,
8 advantages: torch.Tensor,
9 rewards: torch.Tensor,
10 std_reward: float,
11 conf_mask: torch.Tensor,
12 n_conflict: torch.Tensor,
13):
14 mask_f = mask.float(); n_tok = mask_f.sum().clamp(min=1.0)
15
16 idx = input_ids.unsqueeze(-1)
17 new_logp = new_logits.gather(-1, idx).squeeze(-1) - torch.logsumexp(new_logits, dim=-1)
18 ratio = torch.exp(new_logp - new_logp.detach()) # value 1, carries grad
19
20 adv = advantages.squeeze()
21 loss_token = -(ratio * adv * delta)
22 loss_disp = (loss_token * mask_f).sum() / n_tok
23
24 with torch.no_grad():
25 old_logp = old_logits.gather(-1, idx).squeeze(-1) - torch.logsumexp(old_logits, dim=-1)
26 kl = torch.exp(old_logp - new_logp.detach()) - (old_logp - new_logp.detach()) - 1.0
27 mean_kl = (kl * mask_f).sum() / n_tok
28
29 loss = (loss_disp / n_conflict.clamp(min=1)).mean()
30 return loss, n_tok, mean_kl
31
32class UnslothEfficientGTPO(torch.autograd.Function):
33 @staticmethod
34 def forward(
35 ctx,
36 _new_hidden_states: torch.Tensor,
37 _old_hidden_states: torch.Tensor,
38 lm_head: torch.Tensor,
39 comp_ids: torch.Tensor,
40 comp_mask: torch.Tensor,
41 advantages: torch.Tensor,
42 rewards: torch.Tensor,
43 beta: float,
44 scaler=None,
45 n_chunks: int = 1,
46 ):
47 device = _new_hidden_states.device
48 B, G, Lfull, H = _new_hidden_states.shape
49 L = Lfull - 1
50 vocab_size = lm_head.size(0)
51 lm_w = lm_head.t()
52
53 if rewards.dim() == 1:
54 rewards = rewards.unsqueeze(0).expand(B, -1)
55 rewards = rewards.to(device)
56
57 # Compute entropy and KL per completion
58 entropies = torch.zeros((B, G), device=device)
59 for b in range(B):
60 for g in range(G):
61 new_h = _new_hidden_states[b, g]
62 old_h = _old_hidden_states[b, g]
63 ids = comp_ids[b, g]
64 msk = comp_mask[b, g]
65 e, _ = compute_completion_stats(
66 torch.matmul(old_h, lm_w)[:-1], torch.matmul(new_h, lm_w)[:-1], ids, msk
67 )
68 entropies[b, g] = e
69
70 print(f’ENTROPIES : {entropies}’)
71
72 # Adjust advantages using entropy
73 low_ent = entropies <= ENT_THRESHOLD
74
75 # Build delta/conflict masks
76 delta_masks = torch.zeros_like(comp_mask, dtype=torch.float32)
77 conflict_masks = torch.zeros_like(comp_mask, dtype=torch.float32)
78 n_conflict_seq = torch.zeros((B, G), dtype=torch.long, device=device)
79
80 for b in range(B):
81 pos_flags = advantages[b] > 0
82 neg_flags = advantages[b] < 0 # exclude adv == 0
83
84 print(f’POS FLAGS : {pos_flags}’)
85 print(f’NEG FLAGS : {neg_flags}’)
86
87 delta_b, conf_b, nconf_b = build_conflict_mask(
88 comp_ids[b], comp_mask[b], pos_flags, neg_flags, vocab_size
89 )
90
91 active_idx = (advantages[b] != 0).nonzero(as_tuple=True)[0]
92 print(f"Batch {b} -- active conflicting completions: {active_idx.tolist()}")
93
94 for g in range(G):
95 adv_val = advantages[b, g].item()
96 tag = "+" if adv_val > 0 else ("-" if adv_val < 0 else "0")
97 print(f" completion {g}: adv={tag} n_conflict={int(nconf_b[g])}")
98
99 delta_masks[b] = delta_b
100 conflict_masks[b] = conf_b
101 n_conflict_seq[b] = nconf_b
102
103 advantages = torch.where(
104 low_ent,
105 advantages - ENT_SCALE * entropies,
106 torch.zeros_like(advantages)
107 )
108
109 # Prepare for gradient accumulation
110 grad_inputs = torch.empty_like(_new_hidden_states)
111 scaling = scaler.get_scale() if scaler is not None else 1.0
112 acc_loss = torch.zeros(1, device=device)
113 acc_len = torch.zeros(1, device=device)
114 acc_kl = torch.zeros(1, device=device)
115
116 def compute_loss(new_h, old_h, ids, msk, dlt, adv, rew, c_mask, n_conf, scl):
117 new_logits = torch.matmul(new_h, lm_w)[:-1]
118 old_logits = torch.matmul(old_h, lm_w)[:-1]
119 loss_, ln_, kl_ = gtpo_compute_loss(
120 old_logits, new_logits, ids, msk, dlt, beta, adv, rew, std_reward, c_mask, n_conf
121 )
122 return loss_ * scl, (loss_.detach(), ln_, kl_)
123
124 @torch.compile(fullgraph=False)
125 def accumulate(new_h, old_h, ids, msk, dlt, adv, rew, c_mask, n_conf, scl):
126 (g,), (l, (raw, ln, kl)) = torch.func.grad_and_value(compute_loss, argnums=(0,), has_aux=True)(
127 new_h, old_h, ids, msk, dlt, adv, rew, c_mask, n_conf, scl)
128 acc_loss.add_(raw); acc_len.add_(ln); acc_kl.add_(kl)
129 return g
130
131 for b in range(B):
132 for g in range(G):
133 grad_inputs[b, g] = accumulate(
134 _new_hidden_states[b, g], _old_hidden_states[b, g],
135 comp_ids[b, g], comp_mask[b, g], delta_masks[b, g],
136 advantages[b, g].unsqueeze(0), rewards[b, g].unsqueeze(0),
137 conflict_masks[b, g], n_conflict_seq[b, g].unsqueeze(0), scaling
138 )
139
140 ctx.save_for_backward(grad_inputs)
141 grad_inputs_check = grad_inputs.squeeze(0)
142
143 for g in range(grad_inputs_check.shape[0]):
144 grad_norm = grad_inputs_check[g].norm().item()
145 print(f"Completion {g}: grad_norm = {grad_norm:.6f}")
146
147 return acc_loss, acc_len, acc_kl
148
149 @staticmethod
150 def backward(ctx, grad_out, dlen, dkl):
151 (grad_input,) = ctx.saved_tensors
152 return (grad_input,) + (None,) * 12
Listing 3: GTPO Loss